OpenStudy (anonymous):

tan(x−π/2)∗cos(3π/2+x)−sin^3(7π/2−x)/cos(x−π/2)∗tan(3π/2=x)

OpenStudy (anonymous):

plseee solve...

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

yaa i wll give you the options

OpenStudy (anonymous):

a.(1+cos^2x) sin^2x -(1+cos^2x) cos^2x

OpenStudy (anonymous):

@studying? plsee reply

OpenStudy (anonymous):

@Studying? plsee reply

OpenStudy (solomonzelman):

you are familiar with the 2 formulas I hope. 1) \(\large\color{black}{ \bf Tan(c)= \frac{Sin(C)}{Cos(c)} }\) and \(\large\color{black}{ \bf Sin (A+\frac{π}{2}) =Cos(A )}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but then what to do @SolomonZelman

OpenStudy (solomonzelman):

then?

OpenStudy (anonymous):

the solution for the question

OpenStudy (solomonzelman):

You can do it if you also know that Cos(A±180)=Cos(180)Cos(A) ∓ Sin(180)(Sin(A) = Cos(A) ( since Sin(180)= 0 ) And you also know that Cos(A±360)=Cos(360)Cos(A) ∓ Sin(360)(Sin(A) = Cos(A) (since Sin (360) = 0 )

OpenStudy (anonymous):

i know the 1st one but not the 2nd one

OpenStudy (solomonzelman):

You have to understand also that π=180 degree. (You don't need to convert though) You just need to understand the general A and (A+180) and (A+360) and other relations to each other.

OpenStudy (solomonzelman):

The first one, you mean the tan=sin/cos rule ?

OpenStudy (anonymous):

noo..the first means cosa+-360

OpenStudy (solomonzelman):

but do you understand the rules I am posting ?

OpenStudy (anonymous):

yaaa

OpenStudy (solomonzelman):

So what's the problem then? (your score is orange, so I will assume that you are new to this website, but the way it operates, or at least this is the way it should operate; is that the person who answers gives tips and hints (rule or identities) and the ask does the work.)

OpenStudy (anonymous):

see @SolomonZelman i just want you to show me a solution just like you have given in help

OpenStudy (solomonzelman):

First of all, didn't show the solution, that was a hint. Second of all, the asker in that question apparently has no clue about what I said.

OpenStudy (solomonzelman):

I was suspended 3 times and warned a million for posting the solution and doing asker's work. If you really want, you can take it as that I don't know it and just can't do it....